# MoWM: Mixture-of-World-Models for Embodied Planning via Latent-to-Pixel Feature Modulation

### v2

@@ Abstract
Embodied action planning is a core challenge in robotics, requiring models to generate precise actions from visual observations and language instructions. While video generation world models are promising, their reliance on pixel-level reconstruction often introduces visual redundancies that hinder action decoding and generalization. Latent world models offer a compact, motion-aware representation, but overlook the fine-grained details critical for precise manipulation. To overcome these limitations, we propose MoWM, a mixture-of-world-model framework that fuses representations from hybrid world models for embodied action planning. Our approach uses motion-aware representations from a latent model as a high-level prior, which guides the extraction of fine-grained visual features from the pixel-space model. This design allows MoWM to highlight the informative visual details needed for action decoding. Extensive evaluations on the CALVIN benchmark demonstrate that our method achieves state-of-the-art task success rates and superior generalization. We also provide a comprehensive analysis of the strengths of each feature space, offering valuable insights for future research in embodied planning. The code is available at: https://github.com/tsinghua-fib-lab/MoWM.
@@ 1 Introduction
Despite its promise, a key limitation of current world model-based embodied planning lies in the visual representation learning. These models typically rely on features from diffusion-based video generation models (Blattmann et al., 2023; Yang et al., 2024; Wan et al., 2025) pre-trained with pixel-level reconstruction objectives. Such objectives emphasize detailed pixel recovery, yet many robotic tasks do not require perfect reconstruction of all visual elements (Assran et al., 2025). Uniformly encoding all pixels may introduce irrelevant signals that hinder action decoding, complicate the learning of vision-action mappings, and hurt generalization due to overfitting to some task-irrelevant visual details. Alternatively, another line of research explores latent world models (Assran et al., 2025; Zhou et al., 2024; Baldassarre et al., 2025), which learn state transitions in a compressed feature space rather than reconstructing raw pixels. These models employ an encoder to project video sequences into a compact latent representation and a predictor to model dynamics in this space. By focusing on learning motion-aware representations, such methods are more suitable for guiding global action planning. However, they may overlook fine-grained visual details, potentially leading to inaccuracies in tasks requiring dense object interaction.
@@
To address the redundancy of visual representations in video world models, we propose MoWM, a hybrid world model framework for effective embodied planning. Our motivation is to leverage the latent world modeling features as a high-level prior to guide the extraction of fine-grained visual knowledge from the pixel space. This not only eliminates redundant low-level visual information but also avoids the loss of crucial details when using coarse-grained latent features alone. To be specific, our method consists of two stages. In the first stage, we individually train a pixel world model based on video diffusion and a latent world model on embodied manipulation data. Both models are trained to predict future states in their respective spaces conditioned on text instructions. In the second stage, we combine the two world models. The latent world model’s representations are used to modulate the pixel world model’s representations, yielding a fused motion-aware low-level visual representation. In this way, we can direct attention to the visual information most relevant to action decoding, which is then fed into an inverse dynamics model for end-to-end action decoding.
@@
We evaluate our approach on the standard embodied manipulation benchmarks CALVIN (Mees et al., 2022), comparing it against imitation learning-based, VLA-based and world model-based action planning methods. Experimental results demonstrate that our proposed MoWM achieves state-of-the-art performance in task success rates, highlighting the significant potential of hybrid world modeling for embodied action planning. Furthermore, we provide an in-depth analysis of the pixel-level and latent-level visual features during action planning, offering practical guidance for model selection in real-world applications.
The main contributions of this work are summarized as follows:
@@
  • We explored the interactions of visual features from both pixel and latent space world models. We propose an innovative fusion scheme and offer a comprehensive analysis of its effectiveness.
  • Extensive experimental results demonstrate the superiority of our proposed method in both task success rate and generalization to unseen scenarios in embodied action planning.
@@ 3.1 Instruction-conditioned training of hybrid world models
This model is trained on the same embodied datasets, enabling it to anticipate world state transitions from language instructions directly in the latent space.
@@ 4.2 Main Results
Table 2: Ablation study of the latent and pixel feature fusion approaches in MoWM.
@@
MoWM (without fusion) 0.927 0.831 0.741 0.652 0.560 3.70
@@
A key design of our framework is to enhance the feature extraction of the pixel-space world model by incorporating representations from a latent-space world model. To evaluate the effectiveness of this design, we conduct ablation studies comparing three model configurations: (1) MoWM (concat-based fusion): a fusion approach where features from both world models are concatenated and then projected; (2) MoWM (cross attention-based fusion): a fusion approach integrates features from the latent-space world model via cross-attention; (3) MoWM (without fusion): a baseline version that relies solely on the low-level features from the pixel-space world model.
As summarized in Table 2, our ablation study confirms the effectiveness of fusing latent and pixel-space features. The concat-based fusion model achieves the best performance, followed by the modulation-based variant, with the no-fusion baseline performing the worst. These results validate the clear benefit of leveraging latent-space world model features for representation enhancement. The superior performance of the simple concat operation is particularly notable. By preserving the full relevant visual details from low-level features and the motion-aware information from latent features, it creates a more comprehensive visual representation that is better suited for action decoding. Specifically, the concat-based model outperforms the no-fusion variant by an average of 11.2% in task success rate across all five stages, demonstrating the significant value of our hybrid approach. While the attention-based modulation fusion is also beneficial, its performance is lower than simple concatenation. This suggests that while attention can be powerful, it may be challenging to learn the optimal alignment between the two distinct feature spaces, potentially limiting its effectiveness and efficiency compared to a more direct fusion. The no-fusion version performs poorly because the low-level features from the diffusion model are not inherently aligned with action decoding objectives and often contain substantial noise and irrelevant details.
To intuitively demonstrate how the latent world model enhances the pixel-space world model, we conducted a qualitative analysis. As shown in Figure 4, we observed that during action planning, the pixel-space model sometimes produced long periods of static future-state predictions, which is problematic for action decoding. These issues stem from the diffusion model’s training objective, which focuses on fitting all pixels equally, including static background elements, without explicitly emphasizing dynamic motion. This can result in visually plausible but inaccurate motions that negatively impact subsequent action predictions. In contrast, the latent world model learns in a compact space, allowing it to focus more effectively on motion patterns. As illustrated in the examples, this model consistently generates high-quality dynamic patterns and more coherent motions. Consequently, its latent features can guide the low-level pixel features to more accurately reflect the underlying dynamics.
@@ 5 Conclusion and future work
In this paper, we introduced MoWM, a mixture-of-world model framework for embodied action planning that effectively bridges the complementary strengths of pixel-space and latent-space world models. MoWM addresses the key limitation of visual redundancy in pixel-based models by using high-level, motion-aware representations from a latent world model to guide and modulate low-level feature extraction. This design enables the model to suppress task-irrelevant visual details while preserving fine-grained information critical for precise manipulation, resulting in a more insightful visual representation for downstream action decoding. Extensive experiments on the CALVIN benchmark validate the effectiveness of the proposed method.
In the future, we envision several promising directions for further research. A key improvement is to explore dynamic fusion strategies that can adaptively weigh the contributions of latent and pixel features based on task complexity. Furthermore, we will validate our approach on additional benchmarks and real-world robotic platforms to assess its applicability. Finally, we also plan to extend MoWM into a more generalized framework pre-trained on large-scale, unannotated video datasets, enabling zero-shot transfer to a wider range of embodied tasks.

### v3

@@ Abstract
Embodied action planning is a core challenge in robotics, requiring models to generate precise actions from visual observations and language instructions. While video generation world models are promising, their reliance on pixel-level reconstruction often introduces visual redundancies that hinder action decoding and generalization. Latent world models offer a compact, motion-aware representation, but overlook the fine-grained details critical for precise manipulation. To overcome these limitations, we propose MoWM, a mixture-of-world-model framework that fuses representations from hybrid world models for embodied action planning. Our approach combines motion-aware latent world model features with pixel-space features, enabling MoWM to emphasize action-relevant visual details for action decoding. Extensive evaluations on the CALVIN and real-world manipulation tasks demonstrate that our method achieves state-of-the-art task success rates and superior generalization. We also provide a comprehensive analysis of the strengths of each feature space, offering valuable insights for future research in embodied planning. The code is available at: https://github.com/tsinghua-fib-lab/MoWM.
@@ 1 Introduction
Despite their success, a key limitation of current world model–based embodied planning lies in visual representation learning. Most existing methods rely on features from diffusion-based video generation models (Blattmann et al., 2023; Yang et al., 2024; Wan et al., 2025) that operate in pixel space and model dense visual observations. While effective at preserving fine-grained appearance details, such pixel-level representations inherently contain substantial redundant and action-irrelevant information, since all pixels are treated uniformly regardless of their relevance to control. As a result, pixel features often encode large amounts of static background and task-irrelevant content, which can obscure motion cues critical for action decoding and hinder generalization. An alternative line of work explores latent world models (Assran et al., 2025; Zhou et al., 2024; Baldassarre et al., 2025; Karypidis et al., 2024), which learn state transitions in a compressed feature space rather than reconstructing raw pixels. By emphasizing motion-aware and temporally structured representations, these models are more aligned with action planning. However, operating in a highly compressed space may suppress fine-grained visual details, limiting performance in tasks that require precise spatial reasoning or dense object interactions.
@@
To address the redundancy in visual representations learned by pixel-based video world models, we propose MoWM, a hybrid world model framework for effective embodied planning. Our key motivation is to jointly leverage latent and pixel-space representations to construct a visual feature that is both motion-aware and spatially precise. By fusing representations from both spaces, the model can integrate their complementary strengths in a simple yet effective manner. Specifically, MoWM consists of two stages. In the first stage, we independently train a pixel world model based on video diffusion and a latent world model on embodied manipulation data. Both models are trained to predict future states in their respective spaces conditioned on text instructions. In the second stage, we combine the two world models by concatenating their predicted representations at each time step, followed by a lightweight projection to obtain a fused representation that jointly encodes fine-grained visual details and high-level motion dynamics. This hybrid representation preserves the spatial fidelity of pixel-level features while incorporating motion-aware information from the latent space, effectively reducing redundant low-level signals without discarding crucial details. The resulting fused features are then fed into an inverse dynamics model for end-to-end action decoding, enabling the planner to focus on visual information most relevant for control.
@@
We evaluate our approach using the standard embodied manipulation benchmark CALVIN (Mees et al., 2022), comparing it with imitation learning-based, VLA-based, and world model-based action planning methods. Experimental results show that MoWM achieves state-of-the-art task success rates, highlighting the substantial potential of hybrid world modeling for embodied action planning. Additionally, we validate MoWM’s effectiveness on real-world robot manipulation tasks, such as cloth folding. Finally, we provide an in-depth analysis of the pixel-level and latent-level visual features during action planning. Our analysis reveals that while the pixel world model preserves more scene details, it introduces redundant information that hinders action decoding. In contrast, the latent world model captures more dynamic motion, enhancing the action decoding process. The main contributions of this work are summarized as follows:
@@
  • We explored the fusion mechanism of visual features from both pixel and latent space world models, offering a comprehensive analysis of their effectiveness.
  • Extensive experimental results demonstrate the superiority of MoWM in both task success rate and generalization to unseen scenarios in embodied action planning.
@@ 3.1 Instruction-conditioned training of hybrid world models
Both the pixel-space and latent-space world models are trained on the same embodied datasets using an identical temporal sampling strategy. In particular, we align the frame sampling interval and action frequency across the two models, ensuring that each training sample corresponds to the same physical time span and control granularity.
@@ 4.2 Main Results
Figure 4: MoWM successfully executes real-world robotic manipulation tasks, as exemplified by the cloth folding task. This highlights the method’s effectiveness in handling long-horizon, complex tasks in real-world environments.
@@
Real-world robot manipulation task validation. We validate our approach on the real-world AgileX platform. For efficiency, we reduce the controller frequency from 50 Hz to 30 Hz. Policies take as input the robot’s 14 joint angles, camera images, and task descriptions, predicting 30-timestep (1 second) action chunks. The average inference latency is 110 ms for 50 steps. We test the system on a long-horizon cloth folding task, with the instruction ”fold the T-shirt into a rectangle,” which demands high precision in action prediction. As shown in Figure 4, MoWM can successfully execute this challenging real-world robotic manipulation task, demonstrating its effectiveness in handling long-horizon, complex tasks in real-world environments.
@@ 4.3 Ablation Study
Table 2: Ablation study of the latent and pixel feature fusion approaches in MoWM on the CALVIN dataset.
@@
MoWM (only pixel feature) 0.927 0.831 0.741 0.652 0.560 3.70
@@
A key design of our framework is to enhance the feature extraction of the pixel-space world model by incorporating representations from a latent-space world model. To evaluate the effectiveness of this design, we conduct ablation studies comparing three model configurations: (1) MoWM (concat-based fusion): a fusion approach where features from both world models are concatenated and then projected; (2) MoWM (cross attention-based fusion): a fusion approach integrates features from the latent-space world model via cross-attention; (3) MoWM (only pixel feature): a baseline version that relies solely on the low-level features from the pixel-space world model.
As summarized in Table 2, fusing pixel-level and latent representations consistently outperforms the pixel feature-only baseline, validating the benefit of hybrid representations. Among all variants, concat-based fusion achieves the best overall performance, followed by the cross-attention-based fusion, while both no-fusion variants perform significantly worse. Notably, the concat-based model outperforms the no-fusion baseline by an average of 11.2% in task success rate across all five stages, demonstrating the substantial value of integrating latent dynamics into pixel-based representations.
The effectiveness of feature fusion arises from the complementary inductive biases of the two world models. The pixel-space world model preserves fine-grained spatial and appearance details but inherently contains substantial redundancy and noise. As a result, pixel features often encode large amounts of static or irrelevant visual information that can obscure motion cues critical for action decoding, as illustrated in Figure 5. In contrast, the latent-space world model operates in a compact representation space that naturally emphasizes temporal dynamics and motion patterns, yielding features that are more aligned with the requirements of action prediction. This aligns with our observations from feature reconstruction visualizations, where the reconstructed frame sequences faithfully preserve geometric structures and dynamic trajectories but exhibit chromatic deviations, such as altered table colors. This phenomenon implicitly suggests that the latent-space world model prioritizes the encoding of action-centric information while discarding task-irrelevant visual attributes. By injecting latent features into pixel-level representations, the model effectively highlights action-relevant regions while suppressing irrelevant content, producing representations that are spatially precise and dynamically informative.
Although the cross-attention method is expressive, it underperforms simple concatenation in our setting for two possible reasons. First, it introduces additional optimization difficulty by requiring fine-grained alignment between two feature spaces with fundamentally different structures, increasing training sensitivity. Second, pixel features already encode strong spatial structure, making aggressive remapping through attention unnecessary. In contrast, direct concatenation preserves spatial fidelity while injecting motion-aware information in a non-destructive manner, enabling more stable and effective integration.
@@ 5 Conclusion and future work
We introduced MoWM, a hybrid framework for embodied action planning that combines the strengths of pixel-space and latent-space world models. MoWM addresses the visual redundancy in pixel-based models by leveraging motion-aware representations from latent models to enhance low-level feature for action decoding. This approach enables the model to focus on task-relevant details, improving action decoding for precise manipulation. Extensive experiments on the CALVIN benchmark and real-world robot tasks demonstrate the effectiveness of our method.
In the future, we envision several promising directions for further research. A key improvement is to explore dynamic fusion strategies that can adaptively weigh the contributions of latent and pixel features based on task complexity. Furthermore, we also plan to extend MoWM into a more generalized framework pre-trained on large-scale, unannotated video datasets, enabling zero-shot transfer to a wider range of embodied tasks.